OpenStudy (anonymous):

How do I do the Pythagorean Theorem for this? 13 square root 5 + b^2 = 13^2

OpenStudy (nincompoop):

why do you want to use the Pythagorean Theorem?

OpenStudy (anonymous):

Trig Function.

OpenStudy (nincompoop):

? okay… and?

OpenStudy (anonymous):

It says find sec θ if sin θ = 3 square root 13 / 13

OpenStudy (anonymous):

I need to label the triangle of the missing adjacent side.

OpenStudy (nincompoop):

do it.

OpenStudy (anonymous):

I don't know how to do a^2 with the 3 square root 13..

OpenStudy (nincompoop):

you don't know what?

OpenStudy (nincompoop):

you mean cube root?

OpenStudy (anonymous):

\[3\sqrt{13}\]

OpenStudy (nincompoop):

that means \[\sqrt{13\times9}\]

OpenStudy (anonymous):

\[\sqrt{117}\]

OpenStudy (nincompoop):

k what do you do next?

OpenStudy (nincompoop):

are you trying to solve b in 13 square root 5 + b^2 = 13^2 ?

OpenStudy (anonymous):

So \[\sqrt{117}+ b^2 = 169\]

OpenStudy (anonymous):

How do I continue off of that?

OpenStudy (nincompoop):

you might want to isolate b^2 by deducting the whole sqrt(117) on both sides

OpenStudy (anonymous):

Would it just be 169 - \[\sqrt{117}\] on the c side?

OpenStudy (nincompoop):

I don't think your equation was written out properly, but we'll see

OpenStudy (luigi0210):

Pfft, you don't need me, you got the great Nin here

OpenStudy (anonymous):

Sorry I thought he wasn't reading anymore!

OpenStudy (nincompoop):

what c side? now try to get rid of the exponent 2 by obtaining the square root of b^2 and also of the other side of the equation

OpenStudy (anonymous):

\[b^2 = 169 - \sqrt{117}\]

OpenStudy (anonymous):

Is what I'm understanding.

OpenStudy (nincompoop):

just about right

OpenStudy (nincompoop):

if you want to solve for b, you isolate it to one side

OpenStudy (anonymous):

It is isolated.

OpenStudy (nincompoop):

we accomplished that

OpenStudy (nincompoop):

but you have b^2 still a square root of b^2 is b so try to square root both sides of the equation

OpenStudy (anonymous):

b = \[\sqrt{169-\sqrt{117}}\]

OpenStudy (nincompoop):

ya square root of 169 is?

OpenStudy (anonymous):

13.

OpenStudy (nincompoop):

great! now square root of a square root means?

OpenStudy (anonymous):

I'm not sure..

OpenStudy (nincompoop):

it's the same as exponents, you multiply the radical

OpenStudy (anonymous):

So how does that look?

OpenStudy (nincompoop):